40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat May 23 22:43:10 2015, ericq, Update, General, ifo recovery log 
    Reply  Wed May 27 02:53:57 2015, ericq, Update, General, ifo recovery log 
Message ID: 11322     Entry time: Sat May 23 22:43:10 2015     Reply to this: 11326
Author: ericq 
Type: Update 
Category: General 
Subject: ifo recovery log 

Running train-of-thought elog:


East N2 cylinder found empty, replaced. West is >2kpsi


Removed Yuta-specific code from damprestore. A grep for 'yuta' in the python files within /scripts/ shows some other occurances, but nothing that is really in use at this time. New feature of damprestore.py: remembers oplev status.


Ran LSCoffsets. 

WFS offsets relieved (all <20).

Adjusted FSS offset to minimize MC_FAST_MON

ASS ran (but the arm alignment has been astoundingly stable lately. I haven't touched it all this week)

ITMX is the only optic that got a correction over 20 counts. 

BS and *TM oplev spots look well centered, except for ITMX. 

I undid the gain reduction rana introduced because X ASS seemed to be really slow. It is currently fine in its older state. What's going on here?


Some network latency stuff is going on, even freezing up terminals when trying to write text files. This may (or may not) be correlated with the summary page rysnc jobs on nodus. It occurs to me that we have a DAQ ethernet network seperate from the martian network, but the frame transfers need to go through the martian network, since nodus is the only way out to the outside world. If we had a machine/gateway directly from the DAQ network to the caltech network, the martian network wouldn't get bogged down when frames are being uploaded


GTRY = 0.55, ok. Aligned GTRX to 0.52, also ok

Y beatnote was found easily. Have spent >30 minutes looking for X green beatnote. Typical FSS slow and X temperature ranges don't seem to be giving much. Will check the beat alignment with a scope, but if the beat is too high to begin with, it may not work...


I suspect a problem with the X Green BBPD

I could see the IR beatnote between the PSL and AUX X lasers at the input to the frequency counter. (I believe it is a real beatnote because it reacted as expected to the end temperature moving, and stabilizing the end laser to the arm). However, when placing the IR beatnote at a frequency which should've made the green beatnote visible on an analyzer and/or scope, no beatnote was found. I played with the beat alignment to no avail; the DC output of the BBPD behaved as expected, but I never saw anything in the RF output or on the control room analyzer. sad I also checked the beatnote signal chain by hooking up a 1mV 26MHz signal into the cable that hooks up to the RF out of the BBPD; the signal showed up clearly on the control room analyzer. 

I'm not sure what may have happened. ELOG 9996 may be related. 

I'm calling it a night. 

ELOG V3.1.3-